OpenStudy (anonymous):

what is the value of the x variable in the solution to the following system of equations 4x + 2y = 12 x – y = 3

OpenStudy (anonymous):

solve for \(y\) in the second equation, you get \(y=x-3\) substitute that in to the first equation get \[4x+2(x-3)=12\] or \[4x+2x-6=12\] solve that one for \(x\)

OpenStudy (anonymous):

combine like terms to get \[6x-6=12\] then add \(6\) to both sides and divide by 6 to finish

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

yes that would do it

OpenStudy (anonymous):

thank you!